OpenStudy (anonymous):

a 60kg student is standing on a scale in an elevator on Earth. What will be the reading on the scale when the elevator is a) moving down at a constant speed? b) at rest at a floor? c) accelerating at 4.9m/s2 [up] ? d) accerlering at 3.3m/s2 [down]?

OpenStudy (anonymous):

@experimentX ? @mos1635 ?

OpenStudy (anonymous):

i know how to calculate a and b .. just donmt know c and d

OpenStudy (experimentx):

c) mass * (g + accn) d) mass * (g - accn)

OpenStudy (mos1635):

 at any case , reading=N=force from scale at any case total Force=M*a W-N=M*a (if down positive) N-W=M*a (if up positive) give it a shot